OpenStudy (anonymous):

Evaluate the definite integral:

OpenStudy (anonymous):

\[\int\limits_{1}^{4}(\frac{ 1 }{t }+\frac{ 1 }{ t^2 }) dt\]

OpenStudy (anonymous):

$$\int_1^4\left(\frac1t+\frac1{t^2}\right)\,\mathrm{d}x=\left[\log t-\frac1t\right]_1^4=\log4-\frac14-\log1+1=\log4+\frac34$$

OpenStudy (anonymous):

Awesome, thanks!